OpenStudy (anonymous):

[I'm having trouble with part b. how is this part set up on my calculator? Let p=(x,y) be a point on the graph of y=v¯(x) a) Express the distance d to P to the point (1,0) as a function of x My answer: v¯(x^2-x+1) b) Use a graphing utility to graph d=d(x)

OpenStudy (anonymous):

do you know what \[v(x)\] is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what ever it is, the distance will be \[\sqrt{(x-1)^2+v^2(x)}\] and it will be a lot easier to graph and work with if you graph the square of the distance instead, i.e. graph \[(x-1)^2+v^2(x)\]

OpenStudy (anonymous):

if you do not have an expression for \[v(x)\] then there is no way to graph this.

OpenStudy (anonymous):

oh sorry v¯(x^2-x+1) is square root of (x^2-x+1) not v(x)

OpenStudy (anonymous):

i was trying to make a square root sign

OpenStudy (anonymous):

ooooooooooooooooooh i see hold on we can work this out

OpenStudy (anonymous):

the distance between \[(1,0)\] and \[(x,\sqrt{x^2-x+1})\] right?

OpenStudy (anonymous):

messed that up. we get \[d^2=(x-1)^2+x^2-x+1\] \[d^2=x^2-2x+1+x^2-x+1\] \[d^2=2x^2-3x+2\] that looks better

OpenStudy (anonymous):

and as i mentioned it would be easier to work with the square of the distance than the distance, but if you actually need the distance itself, use \[d=\sqrt{2x^2-3x+2}\]

OpenStudy (anonymous):

k